At 3:00 PM a man 135 cm tall casts a shadow 141 cm long. At the same time, a tall building nearby casts a shadow 181 m long. How tall is the building?
Answer by Alan3354(69443)

At 3:00 PM a man 135 cm tall casts a shadow 141 cm long. At the same time, a tall building nearby casts a shadow 181 m long. How tall is the building?
------------
ht/L = ht/L
135/141 = ht/181
ht = 135*181/141 meters